# Vanishing Variance Problem in Fully Decentralized Neural-Network Systems

### v1

@@ Abstract
Federated learning and gossip learning are emerging methodologies designed to mitigate data privacy concerns by retaining training data on client devices and exclusively sharing locally-trained machine learning (ML) models with others. The primary distinction between the two lies in their approach to model aggregation: federated learning employs a centralized parameter server, whereas gossip learning adopts a fully decentralized mechanism, enabling direct model exchanges among nodes. This decentralized nature often positions gossip learning as less efficient compared to federated learning. Both methodologies involve a critical step: computing a representation of received ML models and integrating this representation into the existing model. Conventionally, this representation is derived by averaging the received models, exemplified by the FedAVG algorithm. Our findings suggest that this averaging approach inherently introduces a potential delay in model convergence. We identify the underlying cause and refer to it as the "vanishing variance" problem, where averaging across uncorrelated ML models undermines the optimal variance established by the Xavier weight initialization. Unlike federated learning where the central server ensures model correlation, and unlike traditional gossip learning which circumvents this problem through model partitioning and sampling, our research introduces a variance-corrected model averaging algorithm. This novel algorithm preserves the optimal variance needed during model averaging, irrespective of network topology or non-IID data distributions. Our extensive simulation results demonstrate that our approach enables gossip learning to achieve convergence efficiency comparable to that of federated learning.
[…]
123
1 Introduction
@@
Our investigation of this delay of convergence reveals that Glorot et al. first described such a phenomenon as a "plateau" in single node training [7]. To circumvent this plateau delay, Glorot et al. developed the "Xavier optimal variance" for weight initialization, proposing a methodology to configure model weights accordingly. This weight initialization strategy effectively mitigates the plateau delay in single-node settings, and has become the standard for non-bias weight initialization. In the context of gossip learning, the absence of a central server leads to each node independently initializing their model weights. However, these independently initialized, uncorrelated weight distributions can result in reduced variance following model averaging, thereby compromising back-propagation efficiency and causing the plateau delay.
@@
  • We identify a pervasive phenomenon, termed "plateau delay", in gossip learning systems. Existing studies primarily focus on mitigating this delay rather than eliminating it.
@@
The layout of this paper is as follows: Section 2 provides an overview of federated learning, gossip learning, and Xavier weights initialization. Section 3 discusses the gossip learning system, the plateau delay, methodologies for quantifying plateau delays, and shows how existing federated learning, gossip learning, and transfer learning can circumvent this . Section 4 identifies the vanishing variance as the root cause of the plateau delay. Then, we present our proposed solution, called variance-corrected model averaging, in Section 5. Section 6 shows the performance results of implementing variance-corrected model averaging. Finally, Section 7 summarizes this paper.
2 Background
2.1 Federated Learning
@@
2.2 Gossip Learning
While federated learning offers a novel approach to preserving privacy by keeping data on client devices, it introduces new vulnerabilities such as data poisoning [1], [20], model poisoning [6], privacy attacks [4], [5], and the risk of a single point of failure. In response, some researchers have proposed the removal of the central server to adopt a fully decentralized learning mechanism known as "gossip learning" [9] [13]. Gossip learning eliminates the central server, thereby mitigating the single point of failure and limiting the potential impact of poisoning attacks within a vast peer-to-peer network. Additionally, the increased network distance in gossip learning complicates privacy attacks.
@@ 2.2 Gossip Learning
[…]
@@ 2.3 Xavier Initialization
3 Plateau Delay in Fully Decentralized Neural-Network Systems
3.1 Experiment Setup
This paper starts with the examination of a simple gossip learning architecture, referred to as the "baseline" network. This network consists of 50 nodes, where each node is connected to 8 other peers, forming a regular graph topology with a degree of 8, denoted as $G=Regular(N=50,k=8)$. In this setup, nodes perform a training session by processing a batch of training samples every $T_{acquisition}$ ticks, a unit of virtual time scale within our system. For analytical simplicity, $T_{acquisition}$ is set to 10, making the network operation synchronous, despite gossip learning inherently operating in an asynchronous manner. Upon the completion of training, model weights are sent directly to all connected peers, thus bypassing the Sample and selectPeer steps in Algorithm 2. We combine the Average and Update functions, using the following equation once the model buffer reaches capacity:
@@ 3.1 Experiment Setup
3.2 Plateau Delay in Gossip Learning
@@
Figure 2: Accuracy curves and model weight differences for the baseline network under IID (a,b) and non-IID (c,d) settings. The "plateau delay" is calculated according to Equation 2, and "90% accuracy" marks the point when over 90% of nodes achieve an accuracy higher than 0.9. The "conv1", "conv2", "ip1" and "ip2" are the name of four trainable layers in LeNet.
The accuracy curves presented in Figure 2 reveal a clear convergence delay in the model, which we refer to as the "plateau delay". While the accuracy curve of the LeNet model training on a single node typically exhibits its most rapid increase at the beginning of training, this rapid increase is delayed in our simulations. Thus, we could quantify this delay as the time difference between the start of the training and the most rapid increase, calculated by Equation 2.
@@
3.3 Existing Methods without Plateau Delay
3.3.1 Federated Learning Variant
@@
3.3.2 Weights-Compressed Gossip Learning
@@
3.3.3 Transfer Learning Variant
@@
4 Vanishing Variance Causes Plateau Delay
@@
To overcome this limitation, we investigated an unanswered critical question: why does the plateau delay occur? Through extensive experimentation across various configurations and network topologies, we were able to identify the root cause of the delay. Repeating the experiments described in Section 3.3.1 while introducing a global broadcasting mechanism at the beginning of the model training, whereby an averaged model is calculated and uniformly applied to all nodes, led to the reemergence of the plateau delay. As elaborated in Figure 6, this phenomenon mirrors the delay initially noted in the baseline network as shown in Figure 2.
@@
No plateau delay was noted in the system outlined in Section 3.3.1. Therefore, the plateau delay shown in Figure 6 comes from the global broadcasting mechanism implemented at the beginning of the simulation. Since no training precedes this global broadcasting, this indicates that the averaging process itself triggers the plateau delay. Further examination revealed that Glorot et al. first documented this delay in a single-node training context [7], suggesting that for effective neural network training, initial weights must be set with a specific variance. However, the Average step in Algorithm 2 can reduce this variance in scenarios where the models are not entirely correlated. This reduction in variance for uncorrelated Xavier initialization (both normal and uniform distributions) can be quantified by the following formula:
@@
We define this reduction in variance as the "vanishing variance problem," which also explains how the previously discussed variations of the baseline network effectively counter this problem:
  • Section 3.3.1: the star network architecture, where the central node disseminates its model across the network, ensures the predominance of the central model. This dominance staves off a rapid decrease in variance. However, in scenarios involving global broadcasting (where averaging induces variance reduction) the plateau delay reemerges. Gossip learning implementations that initialize with the same model across nodes similarly mitigate plateau delays by correlating initial models.
  • Section 3.3.2: model compression slows the rate of variance decrease, thereby diminishing the plateau delay.
  • Section 3.3.3: postponing the averaging process until after models have achieved high accuracy helps to preserve variance during training, and preventing the diminishing variance effect of averaging. The strategy of incrementally connecting peers effectively reduces the value of $N$ in Equation 3, further mitigating variance reduction.
@@
5 Variance-Corrected Model Averaging
@@
[…]
@@
[…]
@@
[…]
6 Results
6.1 Comparison with FL and GL
This section shows the convergence efficiency of federated learning and gossip learning frameworks using Algorithm 3 for the Average step. To compare both systems fairly, we matched the number of training nodes: the federated system with 50 clients and one central server, and the gossip system with 50 nodes as per our baseline setup described in Section 3.1. In addition, the averaging weighting factor ($\beta$) in the model averaging algorithm (Equation 1) is set to 0 for gossip learning, to reflect the model replacement strategy in federated learning. However, a non-zero $\beta$ is also considered to account for security concerns within gossip learning environments, where node trustworthiness cannot be guaranteed. Simulations were thus also conducted with $\beta=0.5$, corresponding to the baseline network’s settings.
The comparison of convergence rates between federated learning and gossip learning with variance correction is depicted in Figure 8. For federated learning, we only display the accuracy of the central server. The legend item "First reaching 90% accuracy" denotes the time point for the first node to achieve 0.9 accuracy, while "most reaching 90% accuracy" signifies the time point where over 90% of nodes reach 90% accuracy. The parameter $\alpha=0.5$ in (b)(d)(e) represents the non-IID level, as shown in Figure 1. The figure shows that for both IID (Figure 8(a)(c)) and non-IID (Figure 8(b)(d)) that our variance correction approach improves learning efficiency significantly, allowing gossip learning to catch up with federated learning convergence efficiency.
Figure 8(e) compares the time taken to reach "first reaching 90%" and "most reaching 90%" for gossip learning without variance correction (referred to as "GL" in the legend) and gossip learning with variance correction (referred to as "GL with variance correction"). We could see that applying variance correction could achieve much faster convergence time, nearly 10x faster.
@@ 6.1 Comparison with FL and GL
6.2 Comparison with GL Variants
We compare the performance of variance correction with the combination of weight compression (Section 3.3.2) and the transfer learning (TL) variant (Section 3.3.3). To explore broader scenarios, we modified the following parameters based on the baseline network:
@@ 6.2 Comparison with GL Variants
We simulated all combinations of the above modifications and present the results in Table 1. The left four columns indicate the values of the modified parameters, and the right four columns present the time taken for "most reaching 90%" for the two methods for IID and non-IID data. Our findings suggest that variance correction achieves smaller convergence times than its competitor in all cases. Additionally, variance correction exhibits better scalability, with up to 6x ($T_{acqu}$$=$$10,R$$=$$4,k$$=$$8$) faster convergence compared to its competitor in $N$$=$$1000$ cases. This is because traditional model averaging results in further decrease in variance for larger networks, as indirectly indicated by the increasing $N$ in Equation 3.
Further simulations of gossip learning with only weights compression or only with TL show certain cases where "most reaching 90%" time exceeds 10000 ticks. Due to this poor performance, we do not present these results in Table 1. The cause of such poor performance of weight compression is the increasing network size and insufficient communication caused by compression in non-IID cases. For TL, reduced communication frequency and non-IID data cause nodes to reach high accuracy on their own data making them establish peer connections too rapidly, reducing both accuracy and variance after averaging. This prevents the system from escaping the plateau delay.
Table 1: The time of "most reaching 90%" under various modifications for the two gossip learning systems.

### v2

@@ Abstract
Federated learning and gossip learning are emerging methodologies designed to mitigate data privacy concerns by retaining training data on client devices and exclusively sharing locally-trained machine learning (ML) models with others. The primary distinction between the two lies in their approach to model aggregation: federated learning employs a centralized parameter server, whereas gossip learning adopts a fully decentralized mechanism, enabling direct model exchanges among nodes. This decentralized nature often positions gossip learning as less efficient compared to federated learning. Both methodologies involve a critical step: computing a representation of received ML models and integrating this representation into the existing model. Conventionally, this representation is derived by averaging the received models, exemplified by the FedAVG algorithm. Our findings suggest that this averaging approach inherently introduces a potential delay in model convergence. We identify the underlying cause and refer to it as the ”vanishing variance” problem, where averaging across uncorrelated ML models undermines the optimal variance established by the Xavier weight initialization. Unlike federated learning where the central server ensures model correlation, and unlike traditional gossip learning which circumvents this problem through model partitioning and sampling, our research introduces a variance-corrected model averaging algorithm. This novel algorithm preserves the optimal variance needed during model averaging, irrespective of network topology or non-IID data distributions. Our extensive simulation results demonstrate that our approach enables gossip learning to achieve convergence efficiency comparable to that of federated learning.
I Introduction
@@
Our investigation of this delay of convergence reveals that Glorot et al. first described such a phenomenon as a ”plateau” in single node training [9]. To circumvent this plateau delay, Glorot et al. developed the ”Xavier optimal variance” for weight initialization, proposing a methodology to configure model weights accordingly. This weight initialization strategy effectively mitigates the plateau delay in single-node settings, and has become the standard for non-bias weight initialization. In the context of gossip learning, the absence of a central server leads to each node independently initializing their model weights. However, these independently initialized, uncorrelated weight distributions can result in reduced variance following model averaging, thereby compromising back-propagation efficiency and causing the plateau delay.
@@
  • We identify a pervasive phenomenon, termed ”plateau delay”, in gossip learning systems. Existing studies primarily focus on mitigating this delay rather than eliminating it.
@@
The layout of this paper is as follows: Section II provides an overview of federated learning, gossip learning, and Xavier weights initialization. Section III discusses the gossip learning system, the plateau delay, methodologies for quantifying plateau delays, and shows how existing federated learning, gossip learning, and transfer learning can circumvent this . Section IV identifies the vanishing variance as the root cause of the plateau delay. Then, we present our proposed solution, called variance-corrected model averaging, in Section V. Section VI shows the performance results of implementing variance-corrected model averaging. Finally, Section VII summarizes this paper.
II Background
II-A Federated Learning
@@
II-B Gossip Learning
While federated learning offers a novel approach to preserving privacy by keeping data on client devices, it introduces new vulnerabilities such as data poisoning [10], [11], model poisoning [12], privacy attacks [2], [13], and the risk of a single point of failure. In response, some researchers have proposed the removal of the central server to adopt a fully decentralized learning mechanism known as ”gossip learning” [4] [3]. Gossip learning eliminates the central server, thereby mitigating the single point of failure and limiting the potential impact of poisoning attacks within a vast peer-to-peer network. Additionally, the increased network distance in gossip learning complicates privacy attacks.
@@ II-B Gossip Learning
         $w_{m}\leftarrow\textit{Average}(B)$
         $w\leftarrow\textit{Update}(w,w_{m})$
@@ II-C Xavier Initialization
III Plateau Delay in Fully Decentralized Neural-Network Systems
III-A Experiment Setup
This paper starts with the examination of a simple gossip learning architecture, referred to as the ”baseline” network. This network consists of 50 nodes, where each node is connected to 8 other peers, forming a regular graph topology with a degree of 8, denoted as $G=Regular(N=50,k=8)$. In this setup, nodes perform a training session by processing a batch of training samples every $T_{acquisition}$ ticks, a unit of virtual time scale within our system. For analytical simplicity, $T_{acquisition}$ is set to 10, making the network operation synchronous, despite gossip learning inherently operating in an asynchronous manner. Upon the completion of training, model weights are sent directly to all connected peers, thus bypassing the Sample and selectPeer steps in Algorithm 2. We combine the Average and Update functions, using the following equation once the model buffer reaches capacity:
@@ III-A Experiment Setup
III-B Plateau Delay in Gossip Learning
@@
Figure 2: Accuracy curves and model weight differences for the baseline network under IID (a,b) and non-IID (c,d) settings. The ”plateau delay” is calculated according to Equation 2, and ”90% accuracy” marks the point when over 90% of nodes achieve an accuracy higher than 0.9. The ”conv1”, ”conv2”, ”ip1” and ”ip2” are the name of four trainable layers in LeNet.
The accuracy curves presented in Figure 2 reveal a clear convergence delay in the model, which we refer to as the ”plateau delay”. While the accuracy curve of the LeNet model training on a single node typically exhibits its most rapid increase at the beginning of training, this rapid increase is delayed in our simulations. Thus, we could quantify this delay as the time difference between the start of the training and the most rapid increase, calculated by Equation 2.
@@
III-C Existing Methods without Plateau Delay
III-C1 Federated Learning Variant
@@
III-C2 Weights-Compressed Gossip Learning
@@
III-C3 Transfer Learning Variant
@@
IV Vanishing Variance Causes Plateau Delay
@@
To overcome this limitation, we investigated an unanswered critical question: why does the plateau delay occur? Through extensive experimentation across various configurations and network topologies, we were able to identify the root cause of the delay. Repeating the experiments described in Section III-C1 while introducing a global broadcasting mechanism at the beginning of the model training, whereby an averaged model is calculated and uniformly applied to all nodes, led to the reemergence of the plateau delay. As elaborated in Figure 6, this phenomenon mirrors the delay initially noted in the baseline network as shown in Figure 2.
@@
No plateau delay was noted in the system outlined in Section III-C1. Therefore, the plateau delay shown in Figure 6 comes from the global broadcasting mechanism implemented at the beginning of the simulation. Since no training precedes this global broadcasting, this indicates that the averaging process itself triggers the plateau delay. Further examination revealed that Glorot et al. first documented this delay in a single-node training context [9], suggesting that for effective neural network training, initial weights must be set with a specific variance. However, the Average step in Algorithm 2 can reduce this variance in scenarios where the models are not entirely correlated. This reduction in variance for uncorrelated Xavier initialization (both normal and uniform distributions) can be quantified by the following formula:
@@
We define this reduction in variance as the ”vanishing variance problem,” which also explains how the previously discussed variations of the baseline network effectively counter this problem:
  • Section III-C1: the star network architecture, where the central node disseminates its model across the network, ensures the predominance of the central model. This dominance staves off a rapid decrease in variance. However, in scenarios involving global broadcasting (where averaging induces variance reduction) the plateau delay reemerges. Gossip learning implementations that initialize with the same model across nodes similarly mitigate plateau delays by correlating initial models.
  • Section III-C2: model compression slows the rate of variance decrease, thereby diminishing the plateau delay.
  • Section III-C3: postponing the averaging process until after models have achieved high accuracy helps to preserve variance during training, and preventing the diminishing variance effect of averaging. The strategy of incrementally connecting peers effectively reduces the value of $N$ in Equation 3, further mitigating variance reduction.
@@
V Variance-Corrected Model Averaging
@@
     $\textit{Model}_{avg}\leftarrow Average(B)$triangleright Compute average model
@@
     for Layer $l_{avg}$ in $\textit{Model}_{avg}$ do
5:         for $\textit{Model}_{i}$ in $B$ do
              $\sigma_{li}^{2}\leftarrow\textit{Var}(l_{i})$triangleright Variance of layer $l$ in model $i$
@@
This algorithm ensures that the variance of the averaged model ($\textit{Model}_{avg}$) equals the mean variance of the input models ($\textit{Model}_{i}$). This strategy mitigates the variance reduction identified in Proposition 1 for uncorrelated machine learning models. For correlated models, such as the models recevied by a federated learning server, the impact of the variance rescaling step is minimal, as the ratio $\frac{\sigma_{l}}{\sigma_{lavg}}$ approaches 1, indicating minimal variance deviation.
VI Results
VI-A Comparison with FL and GL
This section shows the convergence efficiency of federated learning and gossip learning frameworks using Algorithm 3 for the Average step. To compare both systems fairly, we matched the number of training nodes: the federated system with 50 clients and one central server, and the gossip system with 50 nodes as per our baseline setup described in Section III-A. In addition, the averaging weighting factor ($\beta$) in the model averaging algorithm (Equation 1) is set to 0 for gossip learning, to reflect the model replacement strategy in federated learning. However, a non-zero $\beta$ is also considered to account for security concerns within gossip learning environments, where node trustworthiness cannot be guaranteed. Simulations were thus also conducted with $\beta=0.5$, corresponding to the baseline network’s settings.
The comparison of convergence rates between federated learning and gossip learning with variance correction is depicted in Figure 8. For federated learning, we only display the accuracy of the central server. The legend item ”First reaching 90% accuracy” denotes the time point for the first node to achieve 0.9 accuracy, while ”most reaching 90% accuracy” signifies the time point where over 90% of nodes reach 90% accuracy. The parameter $\alpha=0.5$ in (b)(d)(e) represents the non-IID level, as shown in Figure 1. The figure shows that for both IID (Figure 8(a)(c)) and non-IID (Figure 8(b)(d)) that our variance correction approach improves learning efficiency significantly, allowing gossip learning to catch up with federated learning convergence efficiency.
Figure 8(e) compares the time taken to reach ”first reaching 90%” and ”most reaching 90%” for gossip learning without variance correction (referred to as ”GL” in the legend) and gossip learning with variance correction (referred to as ”GL with variance correction”). We could see that applying variance correction could achieve much faster convergence time, nearly 10x faster.
@@ VI-A Comparison with FL and GL
VI-B Comparison with GL Variants
We compare the performance of variance correction with the combination of weight compression (Section III-C2) and the transfer learning (TL) variant (Section III-C3). To explore broader scenarios, we modified the following parameters based on the baseline network:
@@ VI-B Comparison with GL Variants
We simulated all combinations of the above modifications and present the results in Table I. The left four columns indicate the values of the modified parameters, and the right four columns present the time taken for ”most reaching 90%” for the two methods for IID and non-IID data. Our findings suggest that variance correction achieves smaller convergence times than its competitor in all cases. Additionally, variance correction exhibits better scalability, with up to 6x ($T_{acqu}$$=$$10,R$$=$$4,k$$=$$8$) faster convergence compared to its competitor in $N$$=$$1000$ cases. This is because traditional model averaging results in further decrease in variance for larger networks, as indirectly indicated by the increasing $N$ in Equation 3.
Further simulations of gossip learning with only weights compression or only with TL show certain cases where ”most reaching 90%” time exceeds 10000 ticks. Due to this poor performance, we do not present these results in Table I. The cause of such poor performance of weight compression is the increasing network size and insufficient communication caused by compression in non-IID cases. For TL, reduced communication frequency and non-IID data cause nodes to reach high accuracy on their own data making them establish peer connections too rapidly, reducing both accuracy and variance after averaging. This prevents the system from escaping the plateau delay.
TABLE I: The time of ”most reaching 90%” under various modifications for the two gossip learning systems.